# Learning to Partially Defer for Sequences

### v1

@@ 6 Experiments
In this section, we demonstrate the quality of a model-based approach to partial deferral in two different experimental setups - News Summarization and Traveling Salesman Problem solver. Code will be made available upon acceptance.
@@ 6.1 Baselines
In experiments, we wish to highlight two aspects of the proposed methodology. The first is the improved cost-accuracy trade-offs in using partial deferrals over whole-sequence deferrals. The second is the advantages of our specific approach to partial deferrals over confidence-based partial deferrals, i.e. thresholding confidence scores from the predictor.
Towards this end, we conduct two experiments. In the first, we compare our token-level and one-time rejectors with a few whole deferral methods. ChowSum, ChowMean, ChowQuantile are thresholding whole-sequence rejectors of different functions of the negative log probability of the generated sequence. WholeModelEmbed is a model-based whole-sequence rejector trained to predict whether the expert is more accurate than the predictor [40].
For the second experiment, our model-based token-level rejector is compared against two confidence-based deferral methods - TokenwiseSoftmax and TokenwiseEntropy. TokenwiseSoftmax rejector defers prediction when the negative log softmax probability of the greedy token exceeds a threshold, and TokenwiseEntropy makes rejection decisions by thresholding the entropy of the predictive distribution. These methods are similar to the confidence-based methods described in Section 2.1.
In the case of one-time deferral, we similarly consider OneTimeSoftmax and OneTimeEntropy as comparative baselines for the one-time deferral method, which choose the token position based on the lowest softmax probability assigned to the greedy token and the highest entropy, respectively, from a set of tokens, ${\cal J}$. Complete descriptions of all baselines can be found in Appendix C.
6.2 Tasks
We refer the reader to Appendix D for training and hyperparameter details of the rejector models.
Traveling Salesman Problem (TSP): The traveling salesman problem is one of the most widely-known NP-Hard optimization problems which requires finding the shortest route to visit every city exactly once and return to the starting city. While no algorithm can efficiently recover the exact solution to this problem, approximate solutions have proven to be useful practically. These methods include the optimal and time-consuming optimization-based surrogates and efficient, yet less optimal ML-based approaches. We, therefore, use our partial deferral method to facilitate collaborations between both approaches.
In this task, the ML predictor is the PointerNet model of [51] trained on randomly generated TSP graphs with $50$ cities (coordinates are drawn from a standard normal distribution), and the optimization-based expert is a Gurobi solver that takes in the partial solution from the predictor as a constraint. The predicted tours are evaluated by the total distance covered by the tour. Since Gurobi solvers do not solve TSP problems in an autoregressive fashion, producing a single city and completing the whole tour will incur similar costs, making the token-level deferral framework unsuitable in this scenario. This, however, highlights a perfect use-case for one-time deferral. $l(y,\widehat{y}_{<j})$ evaluates the total distance of using the subtour $\widehat{y}_{<j}$ and $\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is the sum of the total distance of using the expert-generated subtour or $\widehat{y}^{e}_{\geq j}$ and a cost of $\alpha_{j}$ for deferral that increases monotonically with the number of rejected tokens. More details on $\alpha_{j}$ and its effects on performance can be found in Appendix E.
Extreme Summarization (XSUM): In Extreme Summarization [52], the predictor summarizes long articles in a few sentences. In this case, the predictor is a t5-small model and the expert t5-base model with both models finetuned on the XSUM dataset. All the generated summaries are limited to a token length of $20$ and evaluated using ROUGE1 scores.
[…]
@@ 6.3 Evaluation
[…]
@@ 6.4 Comparison with other deferral methods
Figure 1 and Table 1 show that both model-based token-level deferral and model-based one-time deferral schemes outperform model-based and confidence-based whole deferral schemes with token-level deferral achieving the best performance. Our model-based one-time method also prevails over its counterpart methods and the same holds true for the model-based token-level deferral method. Notably, confidence-based token-level deferral schemes also surpass all other methods in XSUM, demonstrating that token-level deferrals generally offer a superior cost-accuracy tradeoff than whole deferrals. Although model-based one-time deferral exhibits marginal improvement in cost-effectiveness in TSP task, we attribute this to the expert’s lack of ability to predict smaller portions of the sequence at a time. The better cost-accuracy tradeoffs seen in XSUM experiments demonstrate that the advantages of partial deferrals become more prominent with more fine-grained deferrals, which are only possible with access to a suitable expert.
Table 1: Table showing Area under Deferral curve (AUDC) and percentage improvement in AUDC relative to the random rejector for our methods along with the baselines, random, and optimal rejectors. The random rejector makes rejection decisions by flipping a coin with varying probability of landing heads and the optimal rejector achieves the highest possible accuracy regardless of the number of tokens rejected. Numbers in brackets represent the standard deviations measured across the $5$ runs. Low AUDC signifies better cost-accuracy tradeoffs and a positive improvement over a random rejector indicates a shift towards a superlinear tradeoff. The values in bold represent the lowest AUDC and highest percentage improvement for each category (Whole, One-time, Token-level). OneTimeModel scores are reported using the best-performing model trained with ${\cal L}^{\psi}$. See Figure 2 for more details on which performed best.
TSP XSUM
Method AUDC % Improvement AUDC % Improvement
Random 306.12(6.31) 0.00 (0.00) 1433.55 (36.98) 0.00 (0.00)
Optimal 0.00(0.00) 100.00 (0.00) 1212.00 (46.85) 15.48 (1.50)
Whole Deferrals
ChowMean 287.01(7.73) 6.23(2.14) 1430.52 (19.09) 0.13 (3.48)
ChowSum 287.01(7.73) 6.23(2.14) 1431.41 (19.65) 0.06 (3.50)
ChowQuantile 0 301.18(9.40) 1.63(1.18) 1432.28 (25.89) -0.01 (3.89)
ChowQuantile 0.4 287.74(5.91) 6.00(1.16) 1436.10 (24.23) -0.27 (3.80)
ChowQuantile 0.8 288.86(9.35) 5.65(2.12) 1434.48 (19.32) -0.15 (3.40)
ChowQuantile 1.0 295.71(6.12) 3.39(1.42) 1427.27 (18.91) 0.35 (3.50)
WholeModelEmbed 307.88(8.73) -0.58(2.02) 1434.23 (21.57) -0.11 (3.81)
One-time Deferrals
OneTimeModel 270.92(10.67) 11.52(2.41) 1398.39(22.84) 2.35 (4.03)
OneTimeSoftmax 319.79(14.53) -4.44(3.63) 1433.41(26.82) -1.50 (2.13)
OneTimeEntropy 324.23(13.48) -5.93(4.15) 1442.92(24.92) -1.75 (2.38)
Token-level Deferrals
TokenwiseModel - - 1323.92 (25.34) 7.63 (0.93)
TokenwiseSoftmax - - 1354.78 (31.87) 5.48 (0.97)
TokenwiseEntropy - - 1360.41 (31.50) 5.09 (0.63)
Figure 1: Cost-Loss plots for TSP (on the left) and XSUM (on the right) tasks. Deferral curves closer to the lower left corner are more desirable.
@@ 6.5 Size of ${\cal J}$ for One time Deferral
We train one-time rejectors using ${\cal L}^{\psi}$ with varying choices of ${\cal J}$ to study the influence of larger sets of deferral point choices on cost-effectiveness. For both XSUM and TSP, ${\cal J}$ represents a grid of deferral points between $1$ and $L+1$ and we vary the granularity of the grid. Figure 2 demonstrates the need for a modified one-time deferral loss function. The model trained with a ${\cal J}$ including all token positions ($|{\cal J}|=L+1$) is not the most optimal choice with the suboptimality exacerbated in the XSUM experiments. One-time rejectors are generally more cost-effective than its corresponding confidence-based rejectors regardless of the size of ${\cal J}$.
Figure 2: Plots of Size of ${\cal J}$ against percentage improvement in AUDC relative to the random baseline for TSP (on the left) and XSUM (on the right). See Appendix F for exact AUDC values with standard deviations.
@@ Appendix A Proofs for Section 3
It is easy to see that ${\cal L}^{\phi}_{j}(h,r,x,y)$ is convex with respect to $r$ as long as $\phi(\cdot)$ is convex. Since $\phi$ is binary classification calibrated, there is a constant $\gamma$ such that $\gamma\phi(z)\geq\mathbbm{1}_{z\leq 0}$ by Lemma 3 from [53]. Therefore, $\gamma{\cal L}^{\phi}(h,r,x,y)\geq{\cal L}^{\text{Token}}(h,r,x,y)$.
@@
=$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\phi}\left(r.h\right)-{\cal R}^{*}_{\phi}(h)\right)$
@@
Since $\phi$ is binary classification calibrated, we know that ${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\psi^{-1}\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$ where $\psi$ is a non-decreasing convex function, making $\psi^{-1}$ a non-decreasing concave function [53]. By Lemma 3.1, ${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\right)$ where $\tilde{\Gamma}(z)=\left(1+\bar{C}\right)\psi^{-1}\left(\frac{z}{\bar{c}}\right)$. If $\lim_{n\to\infty}{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)=0$, then
@@
=$\overset{(a)}{=}\left(1+\bar{C}\right)\psi\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)}{\bar{c}}\right)$
@@
=$\overset{(b)}{=}0$
(a) by continuity of $\psi$ at 00 and (b) since $\psi(0)=0$ [53]
@@
=${\cal B}=\{(x,y)\mapsto{\cal L}^{\phi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
@@
By the extension of the Talagrand’s Contraction Lemma [54],
@@
where $\tilde{\Gamma}:\mathbb{R}^{+}\to\mathbb{R}^{+}$, ${\cal B}({\cal F})=\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$, and ${\cal A}_{\phi}\left({\cal F},h\right)={\cal R}^{*}_{\phi}\left({\cal F},h\right)-{\cal R}^{*}_{\phi}(h)$ is the approximation error.
@@ Appendix B Proofs for Section 4
${\cal L}^{\psi}(h,r,x,y)$ is a convex upper bound of ${\cal L}^{\text{OneTime}}(h,r,x,y)$ for $\psi\equiv\psi_{\text{ce}}$ and $\psi\equiv\psi_{\text{mae}}$
@@
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+c_{\text{max}}$
If $r(x)\neq j$, then $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq\log 2$ and $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq 0.5$. When $r(x)=j$, $\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq 0$ and $\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq 0$. Therefore, $\gamma_{\text{ce}}\psi_{\text{ce}}({\boldsymbol{g}}(x),j)\geq\mathbbm{1}_{r\left(x\right)\neq j}$ and $\gamma_{\text{mae}}\psi_{\text{mae}}({\boldsymbol{g}}(x),j)\geq\mathbbm{1}_{r\left(x\right)\neq j}$ where $\gamma_{\text{ce}}=\log 0.5$ and $\gamma_{\text{mae}}=2$. So then,
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}+c_{\text{max}}$
$\leq\gamma\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}(x),j)+c_{\text{max}}$
Therefore, $\gamma^{\prime}{\cal L}^{\psi}(h,r,x,y)\geq{\cal L}^{\text{OneTime}}(h,r,x,y)$ as long as $c_{\text{max}}$ is finite
B.3 Proof of Lemma B.3
@@
where ${\cal R}_{\text{multi }0-1}$ is the multiclass 0-1 risk and $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a non-decreasing concave function. If there exists $i,j\in{\cal J}$ such that $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta>0$, then for any ${\boldsymbol{g}}$ over all measurable functions and for any $h$, the following inequality holds for any distribution over ${\cal X}\times{\cal Y}$:
@@
Let $C_{\text{OneTime}}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$ and let $C_{\text{OneTime}}^{*}(h,x)=\inf_{r}C_{\text{OneTime}}(r,h,x)$ be the optimal pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$. Similarly, let $C_{\psi}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}_{\psi}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$ and $C_{\psi}^{*}(h,x)=\inf_{r}C_{\psi}(r,h,x)$ be the optimal pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$. For any measurable function ${\boldsymbol{g}}$
@@
=$\overset{(a)}{=}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$

### v2

@@ 6 Experiments
Our experiments are designed to evaluate two questions: (i) how partial deferral compares to whole-sequence deferral , and (ii) how our model-based approach compares to confidence-based partial deferral. To this end, we consider three tasks (TSP, XSUM, MWP). We refer the reader to Appendix D for training and hyperparameter details of the model rejectors.
6.1 Tasks
Traveling Salesman Problem (TSP): TSP is a classic NP-Hard optimization problems which requires finding the shortest tour visiting all cities exactly once and returning to the start. Optimization-based solvers achieve near-optimality at high computational cost, while ML models are efficient but less accurate. This task highlights a natural collaboration point for partial deferral. We use a PointerNet predictor (Vinyals et al., 2015), trained on synthetic $50$-city TSP graphs (coordinates drawn from a standard normal distribution), and a Gurobi solver as the expert. The expert takes the partial solution from the predictor as a constraint and completes the tour. Tours are evaluated by total distance. Since Gurobi does not operate autoregressively, token-level deferral is unsuitable here; instead, this is a prime use case for one-time deferral. The loss $l(y,\widehat{y}_{<j})$ measures the distance of subtour $\widehat{y}_{<j}$, while the expert cost $\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ combines the distance of the expert-completed subtour $\widehat{y}^{e}_{\geq j}$ with a deferral cost $\alpha_{j}$, which increases monotonically in the number of rejected tokens (see Appendix F).
Extreme Summarization (XSUM). The predictor summarizes long articles in a few sentences using a t5-small model, and the expert is a stronger t5-base model; both are fine-tuned on XSUM (Narayan et al., 2018). We use T5 (encoder–decoder) models because they can condition on any target prefix $\widehat{y}_{<j}$, enabling the expert to complete a partial sequence without exchanging hidden states. By contrast, decoder-only models require reconstructing cross-model KV caches at step $j$, which defeats the purpose of partial feedback. Summaries are limited to 20 tokens and evaluated with ROUGE-1. For token-level deferral, the predictor loss is $l(y,\widehat{y}^{h}_{j})=\bm{1}_{\widehat{y}^{h}_{j}\notin y}$, while the expert cost is $c_{j}(x,\widehat{y}_{<j},y)=\bm{1}_{e(x,\widehat{y}_{<j})\notin y}$. For one-time deferral, we measure $1-\text{ROUGE1}$ score of the expert continuation from step $j$, appended to $\widehat{y}_{<j}$. The joint term $l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ is the sum of this error and the deferral cost $\alpha_{j}$, as in the TSP case.
Multi-Step Weather Prediction (MWP): Given $2$ hours of temperature history ($12$ observations at $10$-minute intervals), the task is to predict the next hour ($L=6$). The predictor is an LSTM trained on real data from the Max Planck Institute in Jena, Germany. The expert is a simulated human: ground-truth values perturbed with zero-mean Gaussian noise. For token-level deferral, the predictor loss is mean-squared error $l(y,\widehat{y}^{h}_{j})=(\widehat{y}^{h}_{j}-y_{j})^{2}$, and the expert cost is $c_{j}(x,\widehat{y}_{<j},y)=(e(x,\widehat{y}_{<j})-y_{j})^{2}$. For one-time deferral, the loss is $l(y,\widehat{y}_{<j})=\|y_{<j}-\widehat{y}_{<j}\|_{2}^{2}$ and $\tilde{c}_{j}(x,\widehat{y}_{<j},y)$ combines this squared error of the expert subsequence $\widehat{y}^{e}_{\geq j}$ and the query cost $\alpha_{j}$.
@@ 6.2 Baselines
To compare the cost-accuracy tradeoffs acheived by partial deferral compared with whole-sequence deferral, we consider three confidence-thresholding baselines—ChowSum, ChowMean, and ChowQuantile—which defer the entire sequence based on the sum, mean, or a chosen quantile of token-level confidence scores computed over the predicted sequence. For TSP and XSUM ($|\mathcal{V}|<\infty$), the per-token score is the negative log–softmax; for MWP ($\mathcal{V}=\mathbb{R}$), it is the Monte Carlo dropout variance. WholeModelEmbed is a model-based whole-sequence rejector trained to predict whether the expert is more accurate than the predictor (Gupta et al., 2024).
For partial deferral, we compare our model-based rejectors against confidence-based methods at both granularities. At the token level, TokenwiseScore defers by thresholding token-level confidence scores, and TokenwiseEntropy defers by thresholding the entropy of the predictive distribution. In the one-time setting, OneTimeScore and OneTimeEntropy select a single handoff position based on the highest token-level confidence and predictive entropy, respectively. Entropy-based baselines apply only to XSUM and TSP, where the output vocabulary $\cal V$} is finite. Formal expressions of these baselines are provided in Appendix C.
@@ 6.3 Evaluation
We evaluate our methods using deferral curves, which plot the system loss against the number of deferred tokens using various thresholds of the rejector score r(x)r(x)(see Figure 1). Loss is task-specific: for TSP, it is the percentage increase in tour length relative to the expert; for XSUM, it is $1-\mathrm{ROUGE1}$; and for MWP, it is the squared error $\|y-\widehat{y}\|_{2}^{2}$. The area under the deferral curve (AUDC) summarizes the cost-quality tradeoff, and we also report the percentage improvement in AUDC over a random rejector, whose trade-off is linear. Thresholds do not directly determine rejection decisions in the one-time setting; to obtain similar curves, we adopt the following policy: if $g_{L+1}(x)$ (the score of the “no deferral” action) exceeds a threshold, the predictor completes the sequence, otherwise we defer at $\arg\max_{j\in\mathcal{J}}g_{j}(x)$. All results are averaged over five runs, each with 100 test instances.
Table 1: Area under the deferral curve (AUDC; lower is better) and percentage improvement relative to a random rejector on TSP, XSUM, and MWP. Random flips a biased coin to defer; Optimal is an oracle that minimizes loss at each deferral budget. Bold indicates the best result (ties bolded) within each block (Whole, One-time, Token-level). Values are mean (std) over five runs (100 test instances per run). OneTimeModel uses the best-performing $\mathcal{L}^{\psi}$ variant (see Figure 2).
TSP XSUM MWP
Method AUDC % Improvement AUDC % Improvement AUDC % Improvement
Random 306.12(6.31) 0.00 (0.00) 1433.55 (36.98) 0.00 (0.00) 555.33 (174.75) 0.00 (0.00)
Optimal 0.00(0.00) 100.00 (0.00) 1212.00 (46.85) 15.48 (1.50) 36.66 (4.34) 93.01 (1.47)
Whole Deferrals
ChowMean 287.01(7.73) 6.23(2.14) 1430.52 (19.09) 0.13 (3.48) 521.75 (151.71) 5.26 (9.92)
ChowSum 287.01(7.73) 6.23(2.14) 1431.41 (19.65) 0.06 (3.50) 521.75 (151.71) 5.26 (9.92)
ChowQuantile 0 301.18(9.40) 1.63(1.18) 1432.28 (25.89) -0.01 (3.89) 528.02 (151.50) 3.91 (9.25)
ChowQuantile 0.4 287.74(5.91) 6.00(1.16) 1436.10 (24.23) -0.27 (3.80) 528.02 (151.50) 3.91 (9.25)
ChowQuantile 0.8 288.86(9.35) 5.65(2.12) 1434.48 (19.32) -0.15 (3.40) 528.02 (151.50) 3.91 (9.25)
ChowQuantile 1.0 295.71(6.12) 3.39(1.42) 1427.27 (18.91) 0.35 (3.50) 528.02 (151.50) 3.91 (9.25)
WholeModelEmbed 307.88(8.73) -0.58(2.02) 1434.23 (21.57) -0.11 (3.81) 576.24 (160.28) -5.28 (15.09)
One-time Deferrals
OneTimeModel 270.92(10.67) 11.52(2.41) 1398.39(22.84) 2.35 (4.03) 508.82 (143.39) 7.23 (11.80)
OneTimeScore 319.79(14.53) -4.44(3.63) 1433.41(26.82) -1.50 (2.13) 521.64 (134.83) 4.46 (10.14)
OneTimeEntropy 324.23(13.48) -5.93(4.15) 1442.92(24.92) -1.75 (2.38) - -
Token-level Deferrals
TokenwiseModel - - 1323.92 (25.34) 7.63 (0.93) 412.45 (131.43) 25.78 (1.33)
TokenwiseScore - - 1354.78 (31.87) 5.48 (0.97) 707.51 (203.38) -28.50 (6.20)
TokenwiseEntropy - - 1360.41 (31.50) 5.09 (0.63) - -
@@ 6.4 Comparison with other deferral methods
Table 1 shows that model-based token-level and one-time deferral schemes outperform whole-sequence baselines (both model- and confidence-based); where applicable, token-level deferral attains the best AUDC. Within each type of granularity, model-based rejectors consistently exceed their confidence-thresholding counterparts. In XSUM, even confidence-based token-level deferral surpasses all whole-sequence methods, highlighting the advantage of fine-grained deferral. Gains for one-time deferral on TSP are modest, which we attribute to the non-autoregressive expert that cannot act on partial subsequences. By contrast, in XSUM and MWP, partial deferral yields markedly better cost–accuracy trade-offs.
Figure 1: Deferral Curve for XSUM (closer to the lower left corner is better). See Appendix E for deferral curves of other tasks.
@@ 6.5 Size of $\cal J$} for One time Deferral
Additonally, we train one-time rejectors with $\mathcal{L}^{\psi}$ with varying candidate sets of deferral points $\mathcal{J}$ to assess how the number of allowable handoff positions affects cost–effectiveness, particularly on XSUM and TSP where the sequence length $L$ is large. For both tasks, $\mathcal{J}$ is a uniform grid over $\{1,\ldots,L{+}1\}$ (with L+1L{+}1 denoting “no deferral”), and we vary the grid granularity. Figure 2 demonstrates the need for a modified one-time deferral loss function. The model trained with all token positions ($|{\cal J}|=L+1$) is not the most optimal choice with the suboptimality exacerbated in the XSUM experiments. One-time rejectors are generally more cost-effective than its corresponding confidence-based rejectors regardless of the size of $\cal J$}.
Figure 2: Plots of Size of $\cal J$} against percentage improvement in AUDC relative to the random baseline for TSP (on the left) and XSUM (on the right). See Appendix G for exact AUDC values.
@@ Appendix A Proofs for Section 3
It is easy to see that ${\cal L}^{\phi}_{j}(h,r,x,y)$ is convex with respect to $r$ as long as $\phi(\cdot)$ is convex. Since $\phi$ is binary classification calibrated, there is a constant $\gamma$ such that $\gamma\phi(z)\geq\mathbbm{1}_{z\leq 0}$ by Lemma 3 from Bartlett et al. [2006]. Therefore, $\gamma{\cal L}^{\phi}(h,r,x,y)\geq{\cal L}^{\text{Token}}(h,r,x,y)$.
@@
$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\phi}\left(r.h\right)-{\cal R}^{*}_{\phi}(h)\right)$
@@
Since $\phi$ is binary classification calibrated, we know that ${\cal R}_{\text{binary }0-1}(f)-{\cal R}^{*}_{\text{binary }0-1}\leq\psi^{-1}\left({\cal R}_{\text{binary }\phi}(f)-{\cal R}^{*}_{\text{binary }\phi}\right)$ where $\psi$ is a non-decreasing convex function, making $\psi^{-1}$ a non-decreasing concave function Bartlett et al. [2006]. By Lemma 3.1, ${\cal R}_{\text{Token}}\left(r,h\right)-{\cal R}_{\text{Token}}^{*}(h)\leq\tilde{\Gamma}\left({\cal R}_{\phi}\left(r,h\right)-{\cal R}^{*}_{\phi}(h)\right)$ where $\tilde{\Gamma}(z)=\left(1+\bar{C}\right)\psi^{-1}\left(\frac{z}{\bar{c}}\right)$. If $\lim_{n\to\infty}{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)=0$, then
@@
$\overset{(a)}{=}\left(1+\bar{C}\right)\psi\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n},h\right)-{\cal R}^{*}_{\phi}(h)}{\bar{c}}\right)$
@@
$\overset{(b)}{=}0$
(a) by continuity of $\psi$ at 0 and (b) since $\psi(0)=0$ Bartlett et al. [2006]
@@
${\cal B}=\{(x,y)\mapsto{\cal L}^{\phi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
@@
By the extension of the Talagrand’s Contraction Lemma Maurer [2016],
@@
where $\tilde{\Gamma}$$\mathbb{R}^{+}\to\mathbb{R}^{+}$, ${\cal B}({\cal F})=\frac{4\rho\sqrt{2\left(\bar{C}^{2}+\bar{l}^{2}\right)}}{L}\sum_{j=1}^{L}\mathfrak{R}_{n}\left({\cal F}_{j}\right)+(\bar{C}+\bar{l})\sqrt{\frac{2\log\left(\frac{2}{\delta}\right)}{n}}$, and ${\cal A}_{\phi}\left({\cal F},h\right)={\cal R}^{*}_{\phi}\left({\cal F},h\right)-{\cal R}^{*}_{\phi}(h)$ is the approximation error.
@@ Appendix B Proofs for Section 4
${\cal L}^{\psi}(h,r,x,y)$ is a convex upper bound of ${\cal L}^{\text{OneTime}}(h,r,x,y)$ for $\psi\equiv\psi_{\text{ce}}$ and $\psi\equiv\psi_{\text{mae}}$ upto some scale $\gamma$ if $c_{\max}>\frac{1}{|\mathcal{J}|-1}\sum_{j\in{\cal J}}\left(l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)$
@@
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $=\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\bm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)$
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\bm{1}_{r\left(x\right)\neq j}$
If $r(x)\neq j$, then $\psi_{\text{ce}}({\bm{g}}(x),j)\geq\log 2$ and $\psi_{\text{mae}}({\bm{g}}(x),j)\geq 0.5$. When $r(x)=j$, $\psi_{\text{ce}}({\bm{g}}(x),j)\geq 0$ and $\psi_{\text{mae}}({\bm{g}}(x),j)\geq 0$. Therefore, $\gamma_{\text{ce}}\psi_{\text{ce}}({\bm{g}}(x),j)\geq\bm{1}_{r\left(x\right)\neq j}$ and $\gamma_{\text{mae}}\psi_{\text{mae}}({\bm{g}}(x),j)\geq\bm{1}_{r\left(x\right)\neq j}$ where $\gamma_{\text{ce}}=\log 0.5$ and $\gamma_{\text{mae}}=2$. So then,
${\cal L}^{\text{OneTime}}\left(h,r,x,y\right)$ $\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\bm{1}_{r\left(x\right)\neq j}$
$\leq\gamma\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\bm{g}}(x),j)$
Therefore, $\gamma{\cal L}^{\psi}(h,r,x,y)\geq{\cal L}^{\text{OneTime}}(h,r,x,y)$
B.2 Proof of Lemma B.3
@@
where ${\cal R}_{\text{multi }0-1}$ is the multiclass 0-1 risk and $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a non-decreasing concave function. If there exists $i,j\in{\cal J}$ such that $\left|l(y,\widehat{y}^{h}_{i})+\tilde{c}_{i}(x,\widehat{y}_{<i},y)-l(y,\widehat{y}^{h}_{j})-\tilde{c}_{i}(x,\widehat{y}_{<i},y)\right|>\Delta>0$, then for any $\bm{g}$} over all measurable functions and for any $h$, the following inequality holds for any distribution over ${\cal X}\times{\cal Y}$:
@@
Let $C_{\text{OneTime}}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$ and let $C_{\text{OneTime}}^{*}(h,x)=\inf_{r}C_{\text{OneTime}}(r,h,x)$ be the optimal pointwise ${\cal L}^{\text{OneTime}}$- risk for a given $x$ and predictor $h$. Similarly, let $C_{\psi}(r,h,x)=\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[{\cal L}^{\text{OneTime}}_{\psi}\left(h,r,x,y\right)\right]$ be the pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$ and $C_{\psi}^{*}(h,x)=\inf_{r}C_{\psi}(r,h,x)$ be the optimal pointwise ${\cal L}^{\psi}$- risk for a given $x$ and predictor $h$. For any measurable function $\bm{g}$}
@@
$\overset{(a)}{=}\mathop{\mathbb{E}}_{y\sim{\cal P}_{Y\mid X=x}}\left[\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\mathbbm{1}_{r\left(x\right)\neq j}-(|{\cal J}|-1)c_{\text{max}}+\sum_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\sum_{j\in{\cal J}}\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right]$